# Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling

### v1

@@ 3.1 UAE-3D: Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
Reconstruction Loss. We define the distance between our reconstructed graph and the original graph $\|\hat{\mathbf{G}}-\mathbf{G}\|$, which includes multiple components to ensure the fidelity of the generated 3D molecules:
@@
where $\mathcal{L}_{\text{atom}}$ and $\mathcal{L}_{\text{bond}}$ are cross-entropy losses for atom and bond types; $\mathcal{L}_{\text{coordinate}}$ measures the MSE between the predicted and the ground truth coordinates; $\mathcal{L}_{\text{distance}}$ is an additionally constraint on the predicted coordinates, to ensure correct inter-atomic distances. The overall reconstruction loss is a weighted sum of these components:
@@
$\mathcal{L}_{\text{UAE-3D}}=\mathcal{L}_{\text{recon}}+\beta\cdot\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\left[D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\mathbf{Z}))\right],$ (21)
@@ 3.2 UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation
Train 1.000 1.000 1.000 0.000 0.585 0.999 0.584 00.251
CDGS 0.991 0.285 0.285 0.285 0.262 0.789 0.022 22.051
JODO 1.000 0.874 0.905 0.902 0.417 0.993 0.483 02.523
MiDi* 0.968 0.633 0.654 0.652 0.392 0.951 0.196 07.054
UDM-3D, ours 1.000 0.879 0.913 0.958 0.525 0.989 0.540 00.692
${}_{\text{3D}}$$\downarrow$ Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
Train 0.861 13.73 1.56E-04 1.81E-04 1.56E-04
EDM 0.831 31.29 4.29E-01 4.96E-01 1.46E-02
JODO 0.845 19.99 8.49E-02 1.15E-02 6.68E-04
MiDi* 0.750 23.14 1.17E-01 9.57E-02 4.46E-03
GeoLDM 0.843 30.68 3.91E-01 4.22E-01 1.69E-02
UDM-3D, ours 0.852 17.36 9.89E-03 5.11E-03 1.74E-04
@@
Train 0.999 0.989 0.989 0.000 0.490 0.992 0.946 0.063
CDGS 0.997 0.951 0.936 0.860* 0.493 0.973 0.784 0.798
JODO 0.999 0.990 0.960 0.780* 0.522 0.986 0.934 0.138
MiDi* 0.998 0.980 0.954 0.769 0.501 0.979 0.882 0.187
UDM-3D, ours 0.999 0.983 0.972 0.948 0.508 0.987 0.897 0.161
${}_{\text{3D}}$$\downarrow$ Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
Train 0.994 0.877 5.44E-04 4.65E-04 1.78E-04
EDM 0.986 1.285 1.30E-01 1.82E-02 6.64E-04
MDM 0.992 4.861 2.74E-01 6.60E-02 2.39E-02
JODO 0.992 0.885 1.48E-01 1.21E-02 6.29E-04
MiDi* 0.983 1.100 8.96E-01 2.08E-02 8.14E-04
GeoLDM 0.989 1.030 2.40E-01 1.00E-02 6.59E-04
UDM-3D, ours 0.993 0.881 7.04E-02 9.84E-03 3.47E-04
@@ 4.4 Training & Sampling Efficiency
4.5 Ablation Study and Hyperparameter Analysis
DiT Backbone. We conduct an ablation study on DiT by comparing it to a vanilla Transformer for de novo 3D molecule generation on the QM9 dataset (Table 5). The results show that DiT outperforms the vanilla Transformer across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, improving diffusion performance.
@@ Appendix A Appendix
A.1 Experimental Details
A.1.1 Evaluation Metrics
@@
A.1.2 Targeted Quantum Properties

### v2

@@ 3.1 UAE-3D: Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
Reconstruction Loss. We define the distance between our reconstructed graph and the original graph $|\hat{\mathbf{G}}-\mathbf{G}|$, which includes multiple components to ensure the fidelity of the generated 3D molecules:
@@
where $\mathcal{L}_{\text{atom}}$ and $\mathcal{L}_{\text{bond}}$ are cross-entropy losses for atom and bond types; $\mathcal{L}_{\text{coordinate}}$ measures the MSE between the predicted and the ground truth coordinates; and $\mathcal{L}_{\text{distance}}$ serves as an additional constraint on the predicted coordinates, to ensure correct inter-atomic distances. In particular, we introduce a bonded distance loss which is incorporated within $\mathcal{L}_{\text{distance}}$. In our implementation, we assign a higher weight to the loss computed over bonded atom pairs to enforce stricter geometric constraints where chemical bonds exist. This weighted scheme ensures that the model captures the intrinsic molecular geometry more accurately and improves the overall VAE reconstruction quality. The overall reconstruction loss is a weighted sum of these components:
@@
$\mathcal{L}_{\text{UAE-3D}}=$ $\mathcal{L}_{\text{recon}}$ (21)
$+$ $\beta\cdot\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\left[D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\mathbf{Z}))\right],$
@@ 3.2 UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation
Train 1.000 1.000 1.000 1.000 0.000 0.585 0.999 0.584 00.251
CDGS 0.991 0.706 0.285 0.285 0.285 0.262 0.789 0.022 22.051
JODO 1.000 0.981 0.874 0.905 0.902 0.417 0.993 0.483 02.523
MiDi* 0.968 0.822 0.633 0.654 0.652 0.392 0.951 0.196 07.054
EQGAT-diff* 1.000 0.989 0.845 0.863 0.859 0.377 0.983 0.161 05.898
UDM-3D, ours 1.000 0.925 0.879 0.913 0.907 0.525 0.990 0.540 00.692
[…]
Train 0.861 0.028 13.73 1.56E-04 1.81E-04 1.56E-04
EDM 0.831 0.002 31.29 4.29E-01 4.96E-01 1.46E-02
JODO 0.845 0.010 19.99 8.49E-02 1.15E-02 6.68E-04
MiDi* 0.750 0.003 23.14 1.17E-01 9.57E-02 4.46E-03
GeoLDM 0.843 0.008 30.68 3.91E-01 4.22E-01 1.69E-02
EQGAT-diff* 0.825 0.007 26.33 1.55E-01 5.21E-02 2.10E-03
UDM-3D, ours 0.852 0.014 17.36 9.89E-03 5.11E-03 1.74E-04
@@
Train 0.999 0.988 0.989 0.989 0.000 0.490 0.992 0.946 0.063
CDGS 0.997 0.951 0.951 0.936 0.860* 0.493 0.973 0.784 0.798
JODO 0.999 0.988 0.990 0.960 0.780* 0.522 0.986 0.934 0.138
MiDi* 0.998 0.976 0.980 0.954 0.769 0.501 0.979 0.882 0.187
EQGAT-diff* 0.999 0.971 0.965 0.950 0.891 0.482 0.950 0.703 2.088
UDM-3D, ours 0.999 0.988 0.983 0.972 0.948 0.508 0.987 0.897 0.161
[…]
Train 0.994 0.953 0.877 5.44E-04 4.65E-04 1.78E-04
EDM 0.986 0.817 1.285 1.30E-01 1.82E-02 6.64E-04
MDM 0.992 0.896 4.861 2.74E-01 6.60E-02 2.39E-02
JODO 0.992 0.934 0.885 1.48E-01 1.21E-02 6.29E-04
GeoLDM 0.989 0.897 1.030 2.40E-01 1.00E-02 6.59E-04
MiDi* 0.983 0.842 1.100 8.96E-01 2.08E-02 8.14E-04
EQGAT-diff* 0.988 0.888 1.520 4.21E-01 1.89E-02 1.24E-03
UDM-3D, ours 0.993 0.929 0.881 7.04E-02 9.84E-03 3.47E-04
@@ 4.1 Experimental Setup
  • GEOM-Drugs Axelrod and Gómez-Bombarelli (2020): Pharmaceutical-scale dataset with 450k drug-like molecules (average 44.4 atoms, max 181 atoms). We use 80%/10%/10% splits and retain the lowest-energy conformer per molecule.
@@ 4.4 Training & Sampling Efficiency
4.5 Ablation Studys
4.5.1 DiT Backbone
We conduct an ablation study on DiT by comparing it to a vanilla Transformer for de novo 3D molecule generation on the QM9 dataset (Table 5). The results show that DiT outperforms the vanilla Transformer across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, improving diffusion performance.
@@ 4.5.1 DiT Backbone
4.5.2 Bonded Distance Loss
Table 6 presents an ablation study on the weight $\gamma_{d}$ for the bonded distance loss on the QM9 dataset. We vary $\gamma_{d}$ to analyze its effect on reconstruction accuracy and the preservation of bond geometries. The results indicate that setting $\gamma_{d}=10$ achieves an optimal balance between compressing coordinate information and maintaining geometric fidelity. This experiment validates the necessity of tuning the bonded distance loss weight to enhance model performance.
Table 6: The influence of different bonded distance loss weight $\gamma_{d}$ on the QM9 dataset.
$\gamma_{d}$ / VAE Rec. Atom Accuracy (%) Bond Accuracy (%) Coordinate RMSD (Å)
0 (w/o $\gamma_{d}$) 1.000 1.000 0.0034
1 1.000 1.000 0.0010
5 1.000 1.000 0.0007
10 1.000 1.000 0.0002
20 1.000 1.000 0.0002
[…]
1 0.996 0.959 0.962 0.940 0.940 0.494 0.978 0.874 0.283
5 0.997 0.965 0.970 0.947 0.947 0.500 0.983 0.888 0.255
10 0.999 0.988 0.983 0.972 0.948 0.508 0.987 0.897 0.161
20 0.999 0.988 0.980 0.969 0.947 0.507 0.987 0.898 0.172
$\gamma_{d}$ / 3D-Metric AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
1 0.985 0.857 1.25E-01 1.80E-02 1.06E-03
5 0.989 0.909 1.12E-01 1.19E-02 7.18E-04
10 0.993 0.929 7.04E-02 9.84E-03 3.47E-04
20 0.993 0.929 7.05E-02 9.84E-03 3.45E-04
4.5.3 SE(3)-Equivariant Augmentations
Table 7 shows the influence of various SE(3)-equivariant augmentations strategies on the performance over the QM9 dataset. The results reveal that data augmentation improves overall performance by providing diverse geometric views during UDM-3D training.
In addition to the standard performance metrics, we further investigate the reconstruction error of UAE-3D under different 3D transformations. Table 8 presents the Root Mean Square Deviation (RMSD) of the reconstructed molecules.
Table 7: The effect of SE(3)-equivariant augmentations strategies on the QM9 dataset.
[…]
w/o aug. 0.995 0.947 0.950 0.935 0.921 0.492 0.971 0.799 0.581
Rot. 0.995 0.948 0.951 0.943 0.927 0.493 0.979 0.875 0.315
Trans. 0.999 0.986 0.980 0.967 0.944 0.507 0.981 0.884 0.202
Trans. + Rot. 0.999 0.988 0.980 0.969 0.947 0.507 0.987 0.898 0.172
3D-Metric AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
w/o aug. 0.993 0.879 1.01E-01 1.18E-02 1.08E-03
Rot. 0.993 0.876 8.12E-02 9.99E-03 3.90E-04
Trans. 0.989 0.896 7.08E-02 1.04E-02 6.81E-04
Trans. + Rot. 0.993 0.929 7.05E-02 9.84E-03 3.45E-04
Table 8: Reconstruction error analysis under different SE(3)-equivariant augmentations strategies on the QM9 dataset.
Augmentation RMSD
w/o aug. 0.0002
Rotation 0.0065
Translation 0.0048
Rotation + Translation 0.0067
4.5.4 Latent Dimensions
To investigate the impact of latent dimensionality on 3D molecule generation, we conduct an ablation study on the QM9 dataset. Table 9 summarizes the performance metrics across different latent space dimensions (4, 8, 16, and 32). The results indicate that a too-low latent dimension (e.g., 4) fails to capture the necessary information, leading to inferior performance. Increasing the dimension to 16 results in substantial performance improvements, while further increasing to 32 does not yield significant additional gains. This demonstrates that a latent dimension of 16 provides an optimal trade-off between model capacity and reconstruction fidelity.
Table 9: The influence of latent space dimensionality for 3D molecule generation on the QM9 dataset.
[…]
4 0.914 0.325 0.580 0.566 0.566 3.27E-01 2.20E-01 4.19E-03
8 0.987 0.882 0.942 0.923 0.923 1.29E-01 1.77E-02 8.14E-04
16 0.999 0.988 0.980 0.969 0.947 7.05E-02 9.84E-03 3.45E-04
32 0.986 0.867 0.943 0.918 0.918 9.98E-02 1.42E-02 7.77E-04
@@ Appendix A Appendix
A.1 Visualization of Representations under Different SE(3) Augmentations
Figure 4 shows the t-SNE visualizations of molecular representations under different SE(3) augmentations. Each dot represents a molecular embedding in the latent space, the colors indicate different molecules, and the color gradient from light to dark corresponds to an increase in the translation distance (left) or the rotation angle (right). Specifically, the left panel (Translation) depicts the embeddings under translation along the same direction, while the right panel (Rotation) illustrates the embeddings under rotation around the same axis.
(a) Translation. Visualization of molecular representations under translation along the same direction. The color gradient indicates increasing translation distance.
(b) Rotation. Visualization of molecular representations under rotation around the same axis with same direction. The color gradient indicates increasing rotation angle.
Figure 4: t-SNE visualizations of molecular representations under different SE(3) augmentations.
A.2 Experimental Details
A.2.1 Evaluation Metrics
@@
A.2.2 Targeted Quantum Properties